OpenStudy (anonymous):

five times a certain odd exceeds its next consecutive integer by 2. find the numbers.

OpenStudy (anonymous):

Is your question right?

OpenStudy (anonymous):

According to me : Let the odd number be \(x\), so its next consecutive will be : \(x+1\).. According to your question: \(5 \times x - (x+1) = 2\) \(5x - x - 1 = 2\) \(4x - 1 = 2\) \[x = \frac{3}{4}\] but this is not right, I am misinterpreting the question I think.. :(

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

May be the next consecutive integer, that is I think saying of odd too.. This way: Let the odd number be \(x\), so its next consecutive will be : \(x+2\).. According to your question: \(5 \times x - (x+2) = 2\) \(5x - x - 2 = 2\) \(4x - 2 = 2\) \(x = 1\)..

OpenStudy (anonymous):

This looks fine to me.. :)

OpenStudy (anonymous):

So, that 'certain odd number' is \(1\) and next consecutive is : \(x+2 = 3..\)

OpenStudy (anonymous):

are you sure ?

OpenStudy (anonymous):

For second one I am sure.. :)

OpenStudy (anonymous):

i have a second question can solve it?

OpenStudy (anonymous):

Unless you interact with me, NO..

OpenStudy (anonymous):

i aggree ur solution..

OpenStudy (anonymous):

12. john can finish their math assignment 4 hours while peter can finish it in 5 hours, how many hours will be required for solving their assignment if the work together ? if peter have been solving their assignment for one hours before john joined him , how many hours will still be required to be able to finish their .